N = 20 alpha = 0.05, 1 tail Null hypothesis: The coin is fair Alternative hypothesis: The coin is biased towards heads Obtained value: 17 heads

a) What is the significance?

b) What is the Critical Value?

c) Do we reject the null?
Answer #1

a) What is the significance?

Significance level = α = 0.05

b) What is the Critical Value?

Critical Z value = 1.645

(by using z-table)

c) Do we reject the null?

Test statistic = Z = (P – p)/sqrt(pq/n)

P = x/n = 17/20

Z = (0.85 - 0.50)/sqrt(0.50*0.50/20) = 3.130495

Z > Critical Z value = 1.645

So, we reject the null hypothesis.